# On a question of Erdős and Gimbel on the cochromatic number

### v1

@@ 1 Introduction
The cochromatic number $\zeta(G)$ of a graph $G$ is the minimum number of colours needed for a vertex colouring where every colour class is either an independent set or a clique. If $\chi(G)$ denotes the usual chromatic number, then clearly $\zeta(G)\leqslant\chi(G)$. Using classical methods, it is not hard to show that for the random graph $G_{n,1/2}$, we have $\zeta({G_{n,1/2}})\sim\chi({G_{n,1/2}})\sim\frac{n}{2\log_{2}n}$.
Erdős and Gimbel [4] (see also [5]) asked the following question: For $G\sim G_{n,1/2}$, does the difference $\chi(G)-\zeta(G)$ tend to infinity as $n\rightarrow\infty$? In other words, is there a function $f(n)\rightarrow\infty$ such that, with high probability [^1] As usual, we say that a sequence $(E_{n})_{n\geqslant 0}$ of events holds with high probability (whp) if ${\mathbb{P}}(E_{n})\rightarrow 1$ as $n\rightarrow\infty$.,
@@ 3 Discussion
If we had a result like Theorem 1, but with the conclusion that $g(n)\geqslant h(n)$ for every $n$ and some $h(n)\gg\sqrt{n}/\log n$, this would imply that the answer to the original question is ‘yes’: by an argument of Alon [2, 9], both $\chi({G_{n,1/2}})$ and $\zeta({G_{n,1/2}})$ are contained in a sequence of intervals of length about $\sqrt{n}/\log n$, respectively, and consequently so is their difference.

### v2

@@ 1 Introduction
The cochromatic number $\zeta(G)$ of a graph $G$ is the minimum number of colours needed for a vertex colouring where every colour class is either an independent set or a clique. If $\chi(G)$ denotes the usual chromatic number, then clearly $\zeta(G)\leqslant\chi(G)$. Using classical methods, it is not hard to show that for the random graph $G_{n,1/2}$, with high probability [^1] As usual, we say that a sequence $(E_{n})_{n\geqslant 0}$ of events holds with high probability (whp) if ${\mathbb{P}}(E_{n})\rightarrow 1$ as $n\rightarrow\infty$. we have $\zeta({G_{n,1/2}})\sim\chi({G_{n,1/2}})\sim\frac{n}{2\log_{2}n}$.
Erdős and Gimbel [4] (see also [5]) asked the following question: For $G\sim G_{n,1/2}$, does the difference $\chi(G)-\zeta(G)$ tend to infinity as $n\rightarrow\infty$? In other words, is there a function $f(n)\rightarrow\infty$ such that, with high probability,
@@
Independently from this note, Raphael Steiner recently also discovered the connection between $\chi(G)-\zeta(G)$ and the concentration interval length of $\chi(G_{n,1/2})$. For his work on this and two other questions of Erdős, Gimbel and Straight, see [10].
@@ 3 Discussion
If we had a result like Theorem 1, but with the conclusion that $g(n)\geqslant h(n)$ for every $n$ and some $h(n)\gg\sqrt{n}/\log n$, this would imply that the answer to the original question is ‘yes’: by an argument of Alon [2, 9], both $\chi({G_{n,1/2}})$ and $\zeta({G_{n,1/2}})$ are contained in a sequence of intervals of length about $\sqrt{n}/\log n$, respectively, and consequently so is their difference. [^3] Formally the statement is: For $G\sim G_{n,1/2}$ and any function $\omega(n)\rightarrow\infty$, there is a sequence of intervals of length at most $\omega(n)\sqrt{n}/\log n$ that contains $\chi(G)-\zeta(G)$ whp. Let $[s_{n},t_{n}]$ be such a sequence of intervals, then taking $g(n)=t_{n}$ would give that $t_{n}\geqslant h(n)\gg\sqrt{n}/\log n$, which implies that $s_{n}\gg\sqrt{n}/\log n$ as well444for a suitable choice of the arbitrary function $\omega(n)\rightarrow\infty$, and so whp $\chi(G)-\zeta(G)\geqslant s_{n}\gg\sqrt{n}/\log n$.